Evaluate   where   

F(x,y,z)=xzi−yzk


  and c is the line segment  from (3,0,1) to (-1,2,0)


1
Expert's answer
2021-05-07T09:06:02-0400

The line segment from "(3,0,1) to (-1,2,0)" is,


"r(t)=\langle 3,0,1\rangle+t\langle -4,2,-1\rangle"


"=\langle 3,0,1\rangle+\langle -4t,2t,-t\rangle"


"=\langle 3-4t,2t,1-t\rangle"


where, "0<t<1"


The line integral is evaluated as,


"\int_{C}F(x,y,z)ds"


"=\int_{C}F(r(t))\cdot r'(t)dt"


"=\int_{0}^{1}\langle (3-4t)(1-t),0,-2t(1-t) \rangle\cdot \langle -4,2,-1 \rangle dt"


"=\int_{0}^{1}(-4(3-7t+4t^2)+2t(1-t)) dt"


"=\int_{0}^{1}(-12+28t-16t^2+2t-2t^2) dt"


"=\int_{0}^{1}(-12+30t-18t^2) dt"


"=\lbrack -12(t)+30(\frac{t^2}{2})-18(\frac{t^3}{3})) \rbrack_{0}^{1}"


"=-12+15-6"


"=-3"


Therefore, the line integral is "\int_{C}F(x,y,z)ds=-3"